# Torsion and Chern-Simons gravity in 4D space-times from a Geometrodynamical four-form

### v2

@@ Abstract
$\ \ \ =\mbox{J}\int{\epsilon^{\mu\nu\alpha\beta}F_{\mu\nu}A_{\alpha}(\ln\chi)_{,\beta}\ d^{4}x}$ (24)24( 24 )
where J is a dimensionless constant and $F_{\mu\nu}=A_{\nu,\mu}-A_{\mu,\nu}$. The action expressed in eq.(24) is invariant under diffeomorphism as well as gauge transformations, $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\xi(x)$.
The upshot of adding $\mathcal{A}_{CS}$ (eq.(24)) to the standard action for an electromagnetic field interacting with charge particles in the presence of gravitation $\mathcal{A}_{EM}$ is that, upon extremizing the full action $\mathcal{A}_{EM}+\mathcal{A}_{CS}$ with respect to $A_{\mu}$, one obtains the following modified Einstein-Maxwell equation,
$F^{\alpha\beta}_{\ \ ;\beta}=-4\pi j^{\alpha}+8\pi\mbox{J}\ w^{\mu\nu\alpha\beta}F_{\mu\nu}\chi_{,\beta}$ (25)25( 25 )
@@
Use of the covariant derivative to the above equation with respect to $x^{\alpha}$ leads to the charge density continuity equation that entails conservation of electric charge $Q$,
$(\sqrt{-g}j^{\alpha})_{,\alpha}=0\ \ \Rightarrow\ \ Q=\int{j^{0}\sqrt{-g}\ d^{3}x}\ \ .$ (26)26( 26 )
Application of the above CS formulation to study its effects on magnetohydrodynamics as well as the $\vec{E}.\vec{B}\neq 0$ regions of a pulsar magnetosphere as a source of propagating $\tilde{w}$ have been studied earlier pdg-ref-47 .
We continue adopting the procedure laid out in the paper by Jackiw and Pi to construct a Chern-Simons action for gravity in the 3+1-dimensional space-time. Instead of their fixed Lorentz vector $v_{\mu}$, we use the covariant derivative of $\phi(x)$ so that,
@@
$\ \ \ =\mbox{H}\int{\epsilon^{\mu\nu\alpha\beta}[\Gamma^{\sigma}_{\nu\tau}\partial_{\alpha}\Gamma^{\tau}_{\beta\sigma}+{2\over{3}}\Gamma^{\sigma}_{\nu\tau}\Gamma^{\tau}_{\alpha\eta}\Gamma^{\eta}_{\beta\sigma}](\ln{\chi})_{,\mu}\ d^{4}x}\ \ \ ,$ (27)27( 27 )
where H is a dimensionless, real constant. After integrating by parts once, eq.(27) can be expressed in terms of the Riemann tensor as,
$\mathcal{A}_{GCS}=-{\mbox{H}\over{2}}\int{\ln{\chi}\ {}^{*}RRd^{4}x}\ \ ,$ (28)28( 28 )
@@
${}^{*}R^{\tau\rho\mu\nu}\equiv{1\over{2}}\epsilon^{\mu\nu\alpha\beta}R^{\tau\rho}_{\ \ \alpha\beta}$ (30)30( 30 )
@@
and $R_{\alpha\beta}=R^{\tau}_{\ \alpha\tau\beta}$. From eq.(27) we note that $\ln{\chi}$ acts like the parameter $\theta$ in Jackiw and Pi ’s paper in which the external vector $v_{\mu}$ is set as $\theta_{,\mu}$. Adding $\mathcal{A}_{CS}+\mathcal{A}_{GCS}$ to $\mathcal{A}$ of eq.(19) and then extremizing the full action with respect to $\phi$ and $g_{\mu\nu}$ leads to,
$\chi^{\ ;\alpha}_{\ ;\alpha}={1\over{2}}\bigg{[}g^{\mu\nu}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi+{\mbox{J}\over{2a_{1}}}\chi\ w^{\mu\nu\alpha\beta}F_{\mu\nu}F_{\alpha\beta}-{\mbox{H}\over{4a_{1}}}\chi\ w^{\mu\nu\alpha\beta}R^{\tau}_{\ \sigma\alpha\beta}R^{\sigma}_{\ \tau\mu\nu}\bigg{]},$ (31)31( 31 )
$R_{\mu\nu}-{1\over{2}}g_{\mu\nu}R={{8\pi}\over{m^{2}_{Pl}}}[T_{\mu\nu}+\Theta_{\mu\nu}+C_{\mu\nu}],$ (32)32( 32 )
@@
$C^{\mu\nu}\equiv-2{\mbox{H}\over{\sqrt{-g}}}\bigg{[}{1\over{4}}{}^{*}RRg^{\mu\nu}-(\ln\chi)_{;\alpha;\beta}\bigg{(}{}^{*}R^{\beta\mu\alpha\nu}+{}^{*}R^{\beta\nu\alpha\mu}\bigg{)}+(\ln\chi)_{,\alpha}\bigg{(}\epsilon^{\alpha\mu\sigma\tau}R^{\nu}_{\ \sigma;\tau}+\epsilon^{\alpha\nu\sigma\tau}R^{\mu}_{\ \sigma;\tau}\bigg{)}\bigg{]}.$ (33)33( 33 )
The first term in the RHS of Eq.(33) is new and is not present in the expression for the Cotton tensor as delineated in Jackiw . It appears in our work because under an infinitesmal variation $g_{\mu\nu}\rightarrow$ $g_{\mu\nu}+\delta g_{\mu\nu}$, the change in $(\ln\chi)_{,\mu}$ occurring in eq.(28) is given by,
$\delta(\ln\chi)_{,\mu}=\delta(\phi_{;\mu}/\phi)=-\delta\Gamma^{\alpha}_{\alpha\mu}=-{1\over{2}}\delta(g^{\alpha\beta}g_{\alpha\beta,\mu})\ \ .$ (34)34( 34 )
However, when the equation of motion for the dynamical four-form given by eq.(31) is substituted in eq.(22), its energy-momentum tensor in the presence of CS-term takes the form,
$\Theta_{\mu\nu}=2a_{1}\bigg{[}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}-{{g_{\mu\nu}}\over{2}}\bigg{(}g^{\alpha\beta}{{\chi_{,\alpha}\chi_{,\beta}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{)}\bigg{]}+{H\over{2\sqrt{-g}}}{}^{*}RRg_{\mu\nu}$ (35)35( 35 )
so that when eq.(35) is substituted in eq.(32) the first term in the RHS of eq.(33) cancels with the last term in the RHS of eq.(35). In other words, the term ${}^{*}RRg_{\mu\nu}$ does not appear in or contribute to the Einstein equations given by eq.(32).
Assuming that the geometrodynamical four-form is a dark energy candidate, the above dynamical equations were studied in an earlier work to address the observed late time acceleration of the expansion rate of the universe (pdg-ref-20 and the references therein). Our objective, in this article, is somewhat different since we propose here that the particle associated with $\tilde{w}$ is a pseudo-scalar which acts as a cold dark matter (CDM) candidate aiding in the formation of supermassive black holes when the universe was young. This is the subject of Section IV.
III.3 (iii) Generalized Exterior Derivative
Apart from the possibility of torsion and Chern-Simons extensions ensuing from the four-form field $\tilde{w}$, there may also be another differential geometric significance of this field. The scalar-density $\phi$ can be used to have an antiderivation acting on antisymmetric tensor-densities of arbitrary weights.
Suppose $\tilde{\alpha}$ is a p-form density with weight $w$ such that its components $\alpha_{\nu_{1}\nu_{2}..\nu_{p}}$ transform to $J^{-w}(x,x^{\prime})\ \alpha_{\nu_{1}\nu_{2}..\nu_{p}}$, under a general coordinate transformation $x\rightarrow x^{\prime}$, $J(x,x^{\prime})$ being the corresponding Jacobian. We define a generalized exterior derivative $\tilde{d}_{w}$ acting on $\tilde{\alpha}$ in the following manner,
@@
$-w\partial_{\mu}(\ln\phi)\ \tilde{d}x^{\mu}\wedge\tilde{\alpha}=\tilde{d}\tilde{\alpha}-w\tilde{d}\ (\ln\phi)\ \wedge\tilde{\alpha}.$
It is easy to see that $\tilde{d}_{w}\tilde{\alpha}$ is a (p+1)-form density of weight $w$. If $\alpha_{1}$ and $\alpha_{2}$ are scalar-densities of weights $w_{1}$ and $w_{2}$, respectively, then the above equation leads to,
$\tilde{d}_{w}\alpha_{i}=\partial_{\mu}\alpha_{i}\ \tilde{d}x^{\mu}-w_{i}\alpha_{i}\ \partial_{\mu}(\ln\phi)\ \tilde{d}x^{\mu},\ \ \ i=1,2$ (36)36( 36 )
which are one-form densities and, furthermore, one can show that,
$\tilde{d}_{w}(\alpha_{1}\tilde{d}_{w}\alpha_{2})=\tilde{d}_{w}\alpha_{1}\wedge\tilde{d}_{w}\alpha_{2}.$ (37)37( 37 )
This generalized exterior derivative also satisfies (a) $\tilde{d}_{w}\tilde{d}_{w}=0$ and (b) $\tilde{d}_{w}(\tilde{\alpha}\wedge\tilde{\beta})=\tilde{d}_{w}\tilde{\alpha}\wedge\tilde{\beta}+(-1)^{p}\tilde{\alpha}\wedge\tilde{d}_{w}\tilde{\beta}$, where $\tilde{\alpha}$ and $\tilde{\beta}$ are p- and q-form densities of weights $w_{1}$ and $w_{2}$, respectively.
These properties are sufficient to qualify $\tilde{d}_{w}$ to be a well defined antiderivation on differential form-densities Mukunda .
From eq.(36) it follows that,
$\tilde{d}_{w}\sqrt{-g}=-\sqrt{-g}\ \tilde{d}\ln({\phi\over{\sqrt{-g}}}),$ (40)40( 40 )
since $\sqrt{-g}$ is a scalar-density of weight +1. There are other physically meaningful tensor densities e.g. the scalar density Dirac delta function and antisymmetric tensor-densities, e.g. dual of $F^{\mu\nu}$, on which $\tilde{d}_{w}$ can act.
In passing, we note that $\tilde{d}_{w}\phi=0$, which is analogous to $g_{\mu\nu;\lambda}=0$.
IV IV. Dark bosons, Bose-Einstein Condensates and Formation of Supermassive Black Holes
The wave-particle duality of quantum mechanics transcends to field-particle duality in quantum field theory. Particles associated with quantum fields are intimately linked with their Poincare’ algebra and special relativistic covariance in the standard framework of quantum field theory (QFT). In the context of the four-form field $\tilde{w}$, or equivalently the scalar-density $\phi(x)$, we may ask what kind of particle is associated with it?
Now, $\phi(x)$ in the geometrically flat Minkowski space-time transforms to,
$\phi\rightarrow\phi^{\prime}=J(t,\vec{r};t,-\vec{r})\ \phi=-\phi\ \ ,$ (41)41( 41 )
@@
Because such a pseudo scalar is likely to interact with matter with a strength at the most comparable with that of gravity, we propose here that $\tilde{w}$ is the candidate for ultra-light cold dark matter (CDM), although such ultralight CDM has been linked with axions in the literature (e.g. see Fukuyama , and the references therein). Can this help us in solving the intriguing problem of supermassive black holes (SMBHs)?
So far, around a million quasars have been observed, each powered by accretion of matter onto SMBHs weighing $\gtrsim 10^{7}M_{\odot}$ pdg-ref-0 ; pdg-ref-1 ; pdg-ref-2 ; Volon . Quasar J0100+2802 found at a redshift, $z=6.33$, is associated with a SMBH of mass $\approx 10^{10}\ M_{\odot}$, while another (J2157–3602) at a redshift, $z=4.7$, has a black hole weighing $\cong 2.4\times 10^{10}\ M_{\odot}$. Similarly, quasars J1120+0641 and J1342+0928 located at redshifts $z=7.09$ and $z=7.54$, respectively, are associated with SMBHs with mass $\gtrsim 10^{8}\ M_{\odot}$. These findings pose a severe challenge to the existing models that invoke seed black holes (BHs) growing via accretion, since one needs extremely large seed BHs with mass $\gtrsim 10^{3}\ M_{\odot}$ at $z\gtrsim 40$ Volon ; pdg-ref-11 .
Recently, using the JWST observations, a very distant SMBH with mass $\sim 10^{6}\ M_{\odot}$ has been discovered at a redshift of 10.6 SMBHatz=10.6 . Of course, even heavier SMBHs at lower redshifts, like with mass $\cong 4\times 10^{10}\ M_{\odot}$ at the centre of Holm 15A galaxy belonging to the galaxy cluster Abell 85 have been detected Abell85 . At a redshift of 3.96, a SMBH weighing $\cong 1.7\times 10^{10}\ M_{\odot}$, accreting matter at a rate one solar mass per day has been seen 17BillionSolarMassBHz=3.962 .
In order to address the formation of SMBHs when the universe was $\lesssim 10^{9}$ yrs old, we had earlier put forward a scenario using the framework of Gross-Pitaevskii equation in which ultra heavy BHs are created because of gravitational contraction of Bose-Einstein condensates (BECs) of ultra-light bosonic CDM PDGEklavya ; PDGFazlu . In this section, we provide a simple semi-classical analysis that captures the essential physics of the problem.
@@
$\lambda_{DB}\sim\frac{h}{p}\gtrsim\bigg{(}\frac{3N}{4\pi R^{3}_{h}}\bigg{)}^{-1/3}=R_{h}\bigg{(}\frac{3N}{4\pi}\bigg{)}^{-1/3}=R_{h}\bigg{(}\frac{3M}{4\pi m}\bigg{)}^{-1/3}$ (42),42( 42 ) ,
where $M$ and $N$ are the total mass of dark bosons and their number, respectively, making up the BEC. Eq.(42) represents the condition for BEC formation since it entails that the de Broglie wavelength be larger than the mean separation between nearby bosons.
In the context of the CDM scenario and assuming a positive cosmological constant, the energy $E_{b}$ of a typical dark boson is given by PDGCosmo ,
$E_{b}\sim\frac{p^{2}}{2m}-\frac{GMm}{R_{h}}+\frac{\Lambda\ r^{2}}{6}mc^{2}<0$ (43).43( 43 ) .
@@
$p^{2}<\frac{2GMm^{2}}{R_{h}}-\frac{\Lambda\ r^{2}}{3}m^{2}c^{2}$ (44).44( 44 ) .
@@
$\Delta p\sim p\gtrsim\frac{\hbar}{2R_{h}}$ (45).45( 45 ) .
From eqs.(42) and (45), we have an inequality,
$\frac{h}{4\pi p}\lesssim R_{h}\lesssim\frac{h}{p}\bigg{(}\frac{3N}{4\pi}\bigg{)}^{1/3}$ (46).46( 46 ) .
@@
Using the result $p\sim\frac{\hbar}{2R_{h}}$ from eq.(45) in eq.(43), we obtain,
$E_{b}(R_{h})\sim\frac{\hbar^{2}}{8mR^{2}_{h}}-\frac{GMm}{R_{h}}+\frac{\Lambda\ R^{2}_{h}}{6}mc^{2}<0$ (47),47( 47 ) ,
@@
$R_{h}\gtrsim\frac{\hbar^{2}}{8GMm^{2}}\bigg{(}1-\frac{R^{3}_{h}\Lambda c^{2}}{6GM}\bigg{)}^{-1}$ (48).48( 48 ) .
@@
Even with a low dark matter halo mass $M\cong 10^{6}\ M_{\odot}$, one finds $\Lambda c^{2}/6GM\cong 10^{-68}\ cm^{-3}$. Hence, for galactic scales $R_{h}\lesssim 10^{22}\ cm$, $R^{3}_{h}\Lambda c^{2}/6GM\ll 1$ so that eq.(48) can be approximated to,
$R_{h}\gtrsim\frac{\hbar^{2}}{8GMm^{2}}\bigg{(}1+\frac{R^{3}_{h}\Lambda c^{2}}{6GM}\bigg{)}\approx\frac{\hbar^{2}}{8GMm^{2}}$ (49).49( 49 ) .
By minimizing $E_{b}$ with respect to $R_{h}$, one can obtain an estimate of the BEC size,
$\frac{\partial E_{b}}{\partial R_{h}}=\frac{\Lambda\ R_{h}}{3}mc^{2}+\frac{GMm}{R^{2}_{h}}-\frac{\hbar^{2}}{4mR^{3}_{h}}=0$ (50)50( 50 )
@@
which leads to a quartic equation to be solved,
$\Lambda R^{4}_{h}+\frac{3GM}{c^{2}}R_{h}-\frac{3\hbar^{2}}{4m^{2}c^{2}}=0$ (51).51( 51 ) .

### v3

@@ 3.2 Chern-Simons extensions
$\ \ \ =\mbox{J}\int{\epsilon^{\mu\nu\alpha\beta}F_{\mu\nu}A_{\alpha}(\ln\chi)_{,\beta}\ d^{4}x}\ ,$ (3.10)
where J is a dimensionless constant and $F_{\mu\nu}=A_{\nu,\mu}-A_{\mu,\nu}$. The action expressed in eq.(3.10) is invariant under diffeomorphism as well as gauge transformations, $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\xi(x)$.
The upshot of adding $\mathcal{A}_{CS}$ to the standard action $\mathcal{A}_{EM}$ for an electromagnetic field interacting with charge particles in the presence of gravitation is that, upon extremising the full action $\mathcal{A}_{EM}+\mathcal{A}_{CS}$ with respect to $A_{\mu}$, one obtains the following modified Einstein-Maxwell equation,
$F^{\alpha\beta}_{\ \ ;\beta}=-4\pi j^{\alpha}+8\pi\mbox{J}\ w^{\mu\nu\alpha\beta}F_{\mu\nu}\ \chi_{,\beta}\ ,$ (3.11)
@@
Use of the covariant derivative to the above equation with respect to $x^{\alpha}$ leads to the standard charge density continuity equation that entails conservation of electric charge $Q$,
$(\sqrt{-g}j^{\alpha})_{,\alpha}=0\ \ \Rightarrow\ \ Q=\int{j^{0}\sqrt{-g}\ d^{3}x}\ \ .$ (3.12)
Application of the above CS formulation to study its effects on magnetohydrodynamics as well as the $\vec{E}.\vec{B}\neq 0$ regions of a pulsar magnetosphere acting as a source of propagating $\tilde{w}$ have been studied earlier [7].
We continue adopting the procedure laid out in the paper by Jackiw and Pi [6] to construct a Chern-Simons action for gravity in the 3+1-dimensional space-time. Instead of their fixed Lorentz vector $v_{\mu}$, we use the covariant derivative of $\phi(x)$ so that,
@@
$\ \ \ =\mbox{H}\int{\epsilon^{\mu\nu\alpha\beta}[\Gamma^{\sigma}_{\nu\tau}\partial_{\alpha}\Gamma^{\tau}_{\beta\sigma}+{2\over{3}}\Gamma^{\sigma}_{\nu\tau}\Gamma^{\tau}_{\alpha\eta}\Gamma^{\eta}_{\beta\sigma}](\ln{\chi})_{,\mu}\ d^{4}x}\ \ \ ,$ (3.13)
where H is a dimensionless, real constant. After integrating by parts once, eq.(3.13) can be expressed in terms of the Riemann tensor as,
$\mathcal{A}_{GCS}=-{\mbox{H}\over{2}}\int{\ln{\chi}\ {}^{*}RR\ d^{4}x}\ \ ,$ (3.14)
@@
${}^{*}R^{\tau\rho\mu\nu}\equiv{1\over{2}}\epsilon^{\mu\nu\alpha\beta}R^{\tau\rho}_{\ \ \alpha\beta}\ .$ (3.16)
@@
and $R_{\alpha\beta}=R^{\tau}_{\ \alpha\tau\beta}$. (We note that $\ln{\chi}$ in eq.(3.13) acts like the parameter $\theta$ in Jackiw and Pi’s paper [6] in which the external vector $v_{\mu}$ is set as $\theta_{,\mu}$.)
Adding $\mathcal{A}_{CS}+\mathcal{A}_{GCS}$ to $\mathcal{A}$ of eq.(3.5) and then extremising the full action with respect to $\phi$ and $g_{\mu\nu}$ leads to,
$\chi^{\ ;\alpha}_{\ ;\alpha}={1\over{2}}\bigg{[}g^{\mu\nu}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi+{\mbox{J}\over{2a_{1}}}\chi\ w^{\mu\nu\alpha\beta}F_{\mu\nu}F_{\alpha\beta}-{\mbox{H}\over{4a_{1}}}\chi\ w^{\mu\nu\alpha\beta}R^{\tau}_{\ \sigma\alpha\beta}R^{\sigma}_{\ \tau\mu\nu}\bigg{]}\ ,$ (3.17)
$R_{\mu\nu}-{1\over{2}}g_{\mu\nu}R={{8\pi}\over{m^{2}_{Pl}}}[T_{\mu\nu}+\Theta_{\mu\nu}+C_{\mu\nu}]\ ,$ (3.18)
@@
$C^{\mu\nu}\equiv-2{\mbox{H}\over{\sqrt{-g}}}\bigg{[}{1\over{4}}{}^{*}RRg^{\mu\nu}-(\ln\chi)_{;\alpha;\beta}\bigg{(}{}^{*}R^{\beta\mu\alpha\nu}+{}^{*}R^{\beta\nu\alpha\mu}\bigg{)}+(\ln\chi)_{,\alpha}\bigg{(}\epsilon^{\alpha\mu\sigma\tau}R^{\nu}_{\ \sigma;\tau}+\epsilon^{\alpha\nu\sigma\tau}R^{\mu}_{\ \sigma;\tau}\bigg{)}\bigg{]}\ .$ (3.19)
The first term in the RHS of Eq.(3.19) is new and is not present in the expression for the Cotton tensor as delineated in [6]. It appears in our work because under an infinitesimal variation $g_{\mu\nu}\rightarrow$ $g_{\mu\nu}+\delta g_{\mu\nu}$, the induced change in $(\ln\chi)_{,\mu}$ occurring in eq.(3.13) (while considering the ensuing variation of $\mathcal{A}_{GCS}$) is given by,
$\delta(\ln\chi)_{,\mu}=\delta(\phi_{;\mu}/\phi)=-\delta\Gamma^{\alpha}_{\alpha\mu}=-{1\over{2}}\delta(g^{\alpha\beta}g_{\alpha\beta,\mu})\ \ .$ (3.20)
However, when the equation of motion for the dynamical four-form given by eq.(3.17) is substituted in eq.(3.8), its energy-momentum tensor in the presence of CS-term takes the form,
$\Theta_{\mu\nu}=2a_{1}\bigg{[}{{\chi_{,\mu}\chi_{,\nu}}\over{\chi}}-{{g_{\mu\nu}}\over{2}}\bigg{(}g^{\alpha\beta}{{\chi_{,\alpha}\chi_{,\beta}}\over{\chi}}+{a_{2}\over{a_{1}}}\chi\bigg{)}\bigg{]}+{H\over{2\sqrt{-g}}}{}^{*}RRg_{\mu\nu}$ (3.21)
so that when eq.(3.21) is substituted in eq.(3.18) the first term in the RHS of eq.(3.19) cancels with the last term in the RHS of eq.(3.21). In other words, the term ${}^{*}RRg_{\mu\nu}$ does not appear in or contribute to the Einstein equations given by eq.(3.18).
Assuming that the geometrodynamical four-form is a dark energy candidate, the above dynamical equations were studied in an earlier work to address the observed late time acceleration of the expansion rate of the universe ([1], and the references therein). Our objective, in this article, is somewhat different since we propose here that the particle associated with $\tilde{w}$ is a pseudo-scalar which acts as a cold dark matter (CDM) candidate aiding in the formation of supermassive black holes when the universe was young. This is the subject of Section IV.
3.3 Generalized Exterior Derivative
Apart from the possibility of torsion and Chern-Simons extensions ensuing from the four-form field $\tilde{w}$, there may be another differential geometric significance of this field. The scalar-density $\phi$ can also be used to have an antiderivation acting on antisymmetric tensor-densities of arbitrary weights.
Suppose $\tilde{\alpha}$ is a p-form density with weight $w$ such that its components $\alpha_{\nu_{1}\nu_{2}..\nu_{p}}$ transform to $J^{-w}(x,x^{\prime})\ \alpha_{\nu_{1}\nu_{2}..\nu_{p}}$, under a general coordinate transformation $x\rightarrow x^{\prime}$, $J(x,x^{\prime})$ being the corresponding Jacobian. We define a generalized exterior derivative $\tilde{d}_{w}$ acting on $\tilde{\alpha}$ in the following manner:
@@ 3.3 Generalized Exterior Derivative
$-w\ \partial_{\mu}(\ln\phi)\ \tilde{d}x^{\mu}\wedge\tilde{\alpha}=\tilde{d}\tilde{\alpha}-w\ \tilde{d}\ (\ln\phi)\ \wedge\tilde{\alpha}\ .$ (3.22)
It is easy to verify that $\tilde{d}_{w}\tilde{\alpha}$ is a (p+1)-form density of weight $w$.
Suppose $\alpha_{1}$ and $\alpha_{2}$ are scalar-densities of weights $w_{1}$ and $w_{2}$, respectively. Then, from eq.(3.22), we have,
$\tilde{d}_{w}\alpha_{i}=\partial_{\mu}\alpha_{i}\ \tilde{d}x^{\mu}-w_{i}\alpha_{i}\ \partial_{\mu}(\ln\phi)\ \tilde{d}x^{\mu},\ \ \ i=1,2$ (3.23)
which are one-form densities and furthermore, one can show that,
$\tilde{d}_{w}(\alpha_{1}\tilde{d}_{w}\alpha_{2})=\tilde{d}_{w}\alpha_{1}\wedge\tilde{d}_{w}\alpha_{2}\ .$ (3.24)
The generalized exterior derivative defined by eq.(3.22) also satisfies (a) $\tilde{d}_{w}\tilde{d}_{w}=0$ and (b) $\tilde{d}_{w}(\tilde{\alpha}\wedge\tilde{\beta})=\tilde{d}_{w}\tilde{\alpha}\wedge\tilde{\beta}+(-1)^{p}\tilde{\alpha}\wedge\tilde{d}_{w}\tilde{\beta}$, where $\tilde{\alpha}$ and $\tilde{\beta}$ are p- and q-form densities of weights $w_{1}$ and $w_{2}$, respectively. These properties are sufficient to qualify $\tilde{d}_{w}$ to be a well defined anti-derivation on differential form-densities [5].
From eq.(3.23) it follows that,
$\tilde{d}_{w}\sqrt{-g}=-\sqrt{-g}\ \tilde{d}\ln({\phi\over{\sqrt{-g}}})\ ,$ (3.25)
since $\sqrt{-g}$ is a scalar-density of weight +1.
There are other physically meaningful tensor densities e.g. Dirac delta function (which is a scalar-density) and anti-symmetric tensor-densities, e.g. $\widetilde{F}^{\mu\nu}=$ dual of $F^{\mu\nu}$, on which $\tilde{d}_{w}$ can act. In passing, we note that $\tilde{d}_{w}\phi=0$, which is analogous to $g_{\mu\nu;\lambda}=0$.
4 Dark bosons, Bose-Einstein Condensates and Formation of Supermassive Black Holes
[…]
Now, $\phi(x)$ in the flat Minkowski space-time transforms to,
$\phi\rightarrow\phi^{\prime}=J(t,\vec{r};t,-\vec{r})\ \phi=-\phi\ \ ,$ (4.1)
@@ 4 Dark bosons, Bose-Einstein Condensates and Formation of Supermassive Black Holes
Because such a pseudo-scalar is likely to interact with matter with a strength at the most comparable with that of gravity, we propose here that $\tilde{w}$ is the candidate for ultra-light cold dark matter (CDM), although such CDM particles have been conventionally linked with axions in the literature (e.g. see [8], and the references therein). Can the dark bosons associated with the geometrodynamical four-form sort out the intriguing problem of the very early formation of supermassive black holes (SMBHs)?
So far, around a million quasars have been observed, each powered by the infall of matter into the deep gravitational potential of SMBHs weighing $\gtrsim 10^{7}M_{\odot}$, and thereby getting overheated to turn into hot plasma [9, 10, 11, 12]. Quasar J0100+2802 found at a redshift, $z=6.33$, is associated with a SMBH of mass $\approx 10^{10}\ M_{\odot}$, while another (J2157–3602) at a redshift, $z=4.7$, has a black hole weighing $\cong 2.4\times 10^{10}\ M_{\odot}$. Similarly, quasars J1120+0641 and J1342+0928 located at redshifts $z=7.09$ and $z=7.54$, respectively, are associated with SMBHs with mass $\gtrsim 10^{8}\ M_{\odot}$. These findings pose a severe challenge to the existing models that invoke seed black holes (BHs) growing via accretion, since one needs extremely large seed BHs with mass $\gtrsim 10^{3}\ M_{\odot}$ at $z\gtrsim 40$ [12, 13].
Recently, using the JWST observations, a very distant SMBH with mass $\sim 10^{6}\ M_{\odot}$ has been discovered at a redshift of 10.6 [14]. Of course, even heavier SMBHs at lower redshifts, like with mass $\cong 4\times 10^{10}\ M_{\odot}$ at the centre of Holm 15A galaxy belonging to the galaxy cluster Abell 85 have been detected [15]. At a redshift of 3.96, a SMBH weighing $\cong 1.7\times 10^{10}\ M_{\odot}$, accreting matter at a rate one solar mass per day has been seen [16].
In order to address the formation of SMBHs when the universe was $\lesssim 10^{9}$ yrs old, we had earlier put forward a scenario using the framework of Gross-Pitaevskii equation in which ultra heavy BHs are created because of gravitational contraction of Bose-Einstein condensates (BECs) of ultra-light bosonic CDM [17, 18]. In this section, we provide a simple semi-classical analysis that captures the essential physics of the problem.
@@
$\lambda_{DB}\sim\frac{h}{p}\gtrsim\bigg{(}\frac{3N}{4\pi R^{3}_{h}}\bigg{)}^{-1/3}=R_{h}\bigg{(}\frac{3N}{4\pi}\bigg{)}^{-1/3}=R_{h}\bigg{(}\frac{3M}{4\pi m}\bigg{)}^{-1/3}\ \ ,$ (4.2)
where $M$ and $N$ are the total mass of dark pseudo-scalars and their number, respectively, making up the BEC. Eq.(4.2) represents the condition for BEC formation as it entails that the de Broglie wavelength of the ultra-light pseudo-scalars be larger than the mean separation between the neighbouring dark bosons.
In the context of a $\Lambda$CDM scenario with a positive cosmological constant present, the energy $E_{b}$ of a typical dark boson is given by [19],
$E_{b}\sim\frac{p^{2}}{2m}-\frac{GMm}{R_{h}}+\frac{\Lambda\ r^{2}}{6}mc^{2}<0\ \ .$ (4.3)
@@
$p^{2}<\frac{2GMm^{2}}{R_{h}}-\frac{\Lambda\ r^{2}}{3}m^{2}c^{2}\ \ .$ (4.4)
@@
$\Delta p\sim p\gtrsim\frac{\hbar}{2R_{h}}\ .$ (4.5)
From eqs.(4.2) and (4.5), we have an inequality,
$\frac{h}{4\pi p}\lesssim R_{h}\lesssim\frac{h}{p}\bigg{(}\frac{3N}{4\pi}\bigg{)}^{1/3}$ (4.6)
@@
Using the result $p\sim\frac{\hbar}{2R_{h}}$ from eq.(4.5) in eq.(4.3), we obtain,
$E_{b}(R_{h})\sim\frac{\hbar^{2}}{8mR^{2}_{h}}-\frac{GMm}{R_{h}}+\frac{\Lambda\ R^{2}_{h}}{6}mc^{2}<0\ \ ,$ (4.7)
@@
$R_{h}\gtrsim\frac{\hbar^{2}}{8GMm^{2}}\bigg{(}1-\frac{R^{3}_{h}\Lambda c^{2}}{6GM}\bigg{)}^{-1}\ \ .$ (4.8)
@@
Even with a low dark matter halo mass $M\cong 10^{6}\ M_{\odot}$, one finds $\lambda_{M}\equiv\Lambda c^{2}/6GM\cong 10^{-68}\ cm^{-3}$. Hence, on galactic scales $R_{h}\lesssim 10^{22}\ cm$, $R^{3}_{h}\Lambda c^{2}/6GM\ll 1$ so that eq.(4.8) can be approximated to,
$R_{h}\gtrsim\frac{\hbar^{2}}{8GMm^{2}}\bigg{(}1+\frac{R^{3}_{h}\Lambda c^{2}}{6GM}\bigg{)}\approx\frac{\hbar^{2}}{8GMm^{2}}\ \ .$ (4.9)
In a BEC, most of the identical bosons occupy the ground state. Therefore, minimizing $E_{b}$ (given by eq.(4.7)) with respect to $R_{h}$, one obtains,
$\frac{\partial E_{b}}{\partial R_{h}}=\frac{\Lambda\ R_{h}}{3}mc^{2}+\frac{GMm}{R^{2}_{h}}-\frac{\hbar^{2}}{4mR^{3}_{h}}=0\ ,$ (4.10)
@@
which leads to a quartic equation that needs to be solved in order to estimate the BEC size,
$\Lambda R^{4}_{h}+\frac{3GM}{c^{2}}R_{h}-\frac{3\hbar^{2}}{4m^{2}c^{2}}=0\ \ .$ (4.11)